find the molarity of the solution in which there is 60grams of sodium hydroxide in two liters of solution.
1
Expert's answer
2013-06-18T10:32:49-0400
At first we should calculate the amount of substance (n) of sodium hydroxide:
n(subs.) = m(subs.)/M(subs.)
M(NaOH) = 23 + 16 + 1 = 40 g/mol
n(NaOH) = 60g/(40 g/mol) = 1,5 mol
We can obtain molarity (c) using formula:
c(substance) = n(substance)/V(solution)
c(NaOH) = 1,5 mol / 2 l = 0,75 mol∙l-1